OpenStudy (anonymous):

How do I do this step by step... x - 31 - (1.3x - 31)

OpenStudy (anonymous):

That just disappeared...

jigglypuff314 (jigglypuff314):

x - 31 - (1.3x - 31) distribute the negative x - 31 - 1.3x + 31 rearrange with Commutative Property x - 1.3x - 31 + 31 identity property x - 1.3x + 0 add like terms 1x - 1.3x (or distribute/factor) (1 - 1.3)x subtract ...

sammixboo (sammixboo):

JUST SAYING!"1.3" was replaced by "13/10". 13x/10 - 31 31 = 31/10 = 31 • 10/10 13x - (31 • 10)/10 = 13x - 310/10 x - 31 - 13x-310/10 x - 31 = x - 31/1 = (x - 31) • 10/10 (x-31) • 10 - ((13x-310))/10 =-3x/10 -3x/10

sammixboo (sammixboo):

Yeah I saw that too @wallenbaughz